OpenStudy (anonymous):

My numerator is 2 less than my denominator. I am more than 1/2. What fraction can i be?

OpenStudy (one098):

Do you have a guess?

OpenStudy (one098):

It's okay. I'll help you.

OpenStudy (anonymous):

yay . thanks

OpenStudy (one098):

We're not going to make this complicated when its not. If its a fraction more then 1/2 and the denominator is 2 more then the numerator, cant we multiply 1/2 by 2 to get the other fraction?

OpenStudy (one098):

It qualifies. It's more then 1/2 and the denominator would be 2 more then the numerator.

OpenStudy (one098):

Do you need more help?

OpenStudy (anonymous):

so you just multiply 1/2 by 2 and that's the answer ?

OpenStudy (one098):

Yes. So what do you get?

OpenStudy (anonymous):

it's 1 right ?

OpenStudy (one098):

:/ \[\frac{ 1 }{ 2 } \times \frac{ 2 }{ 2 }\] What do you get?

OpenStudy (anonymous):

it's 1/2 .. so the answer is 1/2 ? wow .. ! so that's it //.. thanks @One098

OpenStudy (one098):

If you multiply 1 by 2, you get 2. If you multiply 2 by 2, you get 4. So its 2/4.

OpenStudy (one098):

Numerator is 2 less then the denominator.

OpenStudy (anonymous):

ahh . THANK YOUU SO MUCH for helping :D

OpenStudy (one098):

You're welcome, anytime!